# CTI-REALM : Benchmark to Evaluate Agent Performance on Security Detection Rule Generation Capabilities

### 13 Mar 2026 v1

@@ 4 Experimental Setup
Models Evaluated: We assess 16 model configurations: three Anthropic Claude models (Opus 4.6 with high reasoning effort [22], Opus 4.5, Sonnet 4.5), nine OpenAI GPT-5 family variants (GPT-5, 5.1, 5.2 each at high/medium/low reasoning effort), GPT-5-Mini, GPT-4.1, and two reasoning models (O3, O4-Mini). The reasoning effort levels control the computational budget for extended thinking, enabling analysis of how reasoning depth affects detection performance.
@@ 5.1 Overall Model Performance
Claude Opus 4.6 (High) Anthropic 0.6373 0.0374 [0.562, 0.712] 0.266 0.372 0.68 0.58 0.94 0.86 31
Claude Opus 4.5 Anthropic 0.6244 0.0338 [0.556, 0.692] 0.287 0.337 0.74 0.66 1.00 0.92 32
Claude Sonnet 4.5 Anthropic 0.5872 0.0328 [0.521, 0.653] 0.278 0.310 0.70 0.62 1.00 0.88 37
GPT-5 (Med) OpenAI 0.5720 0.0337 [0.504, 0.640] 0.235 0.337 0.55 0.64 1.00 0.38 32
GPT-5.2 (Med) OpenAI 0.5716 0.0325 [0.506, 0.637] 0.226 0.346 0.47 0.46 0.98 0.72 31
GPT-5 (High) OpenAI 0.5633 0.0329 [0.497, 0.629] 0.230 0.333 0.54 0.64 0.96 0.40 32
GPT-5.2 (High) OpenAI 0.5513 0.0367 [0.478, 0.625] 0.232 0.319 0.47 0.50 0.94 0.84 35
GPT-5 (Low) OpenAI 0.5413 0.0315 [0.478, 0.605] 0.204 0.337 0.45 0.54 0.98 0.22 25
GPT-5.1 (Med) OpenAI 0.5133 0.0300 [0.453, 0.574] 0.244 0.269 0.55 0.72 1.00 0.42 26
GPT-5.1 (High) OpenAI 0.4946 0.0373 [0.420, 0.570] 0.238 0.256 0.55 0.66 0.92 0.56 29
GPT-5.2 (Low) OpenAI 0.4898 0.0248 [0.440, 0.540] 0.221 0.269 0.41 0.62 1.00 0.50 27
O3 OpenAI 0.4707 0.0281 [0.414, 0.527] 0.204 0.267 0.40 0.42 0.92 0.68 33
GPT-5-Mini OpenAI 0.4506 0.0292 [0.392, 0.509] 0.188 0.263 0.51 0.22 0.98 0.20 27
GPT-4.1 OpenAI 0.4186 0.0235 [0.371, 0.466] 0.193 0.225 0.55 0.32 1.00 0.02 21
GPT-5.1 (Low) OpenAI 0.3731 0.0211 [0.331, 0.415] 0.194 0.179 0.49 0.42 1.00 0.06 22
O4-Mini OpenAI 0.3602 0.0238 [0.312, 0.408] 0.167 0.193 0.38 0.26 0.94 0.20 26
The results reveal several notable patterns across the 0.637–0.419 performance range. Anthropic Claude models occupy the top three positions, with a clear separation from the best OpenAI result (GPT-5 Med, 0.572). This advantage appears driven by stronger tool-use and agentic capabilities: Claude models achieve substantially higher C3 (query execution) rewards (0.86–0.92 vs. below 0.50 for most OpenAI models) and more consistent CTI comprehension at C0 (0.68–0.74 vs. 0.38–0.55), suggesting more effective iterative investigation behavior.
@@ 5.3 Checkpoint Analysis
Figure 4(b) shows normalized mean rewards for each checkpoint. C3 (query execution) is the most discriminating - Claude models score 0.86–0.92 while most OpenAI models fall below 0.50, with GPT-4.1 at just 0.02. C1 (MITRE mapping) varies widely (0.22–0.72), with GPT-5.1 Med leading at 0.72 despite a modest overall reward, suggesting strong CTI knowledge does not guarantee downstream performance. C0 (CTI report analysis) separates Claude (0.68–0.74) from most OpenAI models (0.38–0.55), reflecting differences in threat intelligence comprehension that propagate through subsequent stages.
@@ 5.6 Ablation: Impact of CTI Tools
Table 3: Ablation study: impact of removing CTI-specific tools. $\Delta$ = Full - Minimal.
Model Full Score Minimal Score $\Delta$ Cohen’s $d$
Claude Opus 4.6 (High) 0.6130 0.5358 +0.0772 0.30
Claude Opus 4.5 0.6124 0.4624 +0.1500 0.59
GPT-5 (Med) 0.5454 0.4120 +0.1335 0.64
GPT-5.2 (Med) 0.5331 0.4555 +0.0776 0.35
GPT-5.1 (Med) 0.4815 0.3641 +0.1174 0.61
All five models show statistically meaningful performance degradation without CTI tools. Claude Opus 4.5 shows the largest absolute drop ($\Delta=-0.150$), falling from 0.612 to 0.462. GPT-5 (Med) drops by 0.134, and GPT-5.1 (Med) by 0.117. Even the top model, Claude Opus 4.6 (High), drops by 0.077.
Since removing CTI tools mechanically zeroes C0, we exclude it from the per-checkpoint comparison. The right side of Table 3 shows the remaining deltas (positive = full toolkit better). C1 (MITRE mapping) is negative for all models, indicating agents compensate by calling search_mitre_techniques more aggressively without CTI context. C3 (query execution) is positive for three of five models, suggesting agents struggle to construct effective queries without CTI guidance. C4 (detection quality) is positive across all models (+0.042 to +0.445), confirming that the overall reward drop is driven by degraded detection rule quality. CTI context is essential for producing high-quality detections, even if intermediate checkpoints can be reached through alternative paths.
@@ 5.7 Memory Augmentation: Closing the Gap with Larger Models
Memory augmentation closes 33% of the 0.184-point gap to GPT-5 (Med). The improvement is concentrated in knowledge-dependent checkpoints: C1 (MITRE mapping) doubles from 0.210 to 0.420, and the ground truth reward improves by 19%. However, C3 (query execution) remains unchanged at 0.150, suggesting query construction is an inherent model capability rather than a knowledge gap. The remaining gap to GPT-5 lies primarily in C3 (GPT-5 Med achieves 0.500) and detection quality, indicating the larger model’s advantage is in iterative query refinement rather than domain knowledge.
@@
Model Score Checkpoint Reward Ground Truth Reward C0 C1 C2 C3 Steps
GPT-5 (Med) 0.5556 2.295 3.261 0.589 0.510 1.000 0.200 34
GPT-5 (Low) 0.5478 2.166 3.312 0.586 0.480 1.000 0.100 27
GPT-5 (High) 0.5348 2.487 2.861 0.672 0.510 1.000 0.320 37
GPT-5-Mini (Memory) 0.4324 2.071 2.253 0.602 0.420 1.000 0.060 41
GPT-5-Mini 0.3714 1.822 1.892 0.607 0.210 0.960 0.060 27
@@ 8 Discussion and Conclusion
We introduced CTI-REALM, a benchmark for evaluating AI agents on realistic detection engineering workflows. Our evaluation of 16 frontier models reveals that medium reasoning effort consistently outperforms high and low settings. Anthropic Claude models lead overall (Opus 4.6 High: 0.637), and CTI tools provide genuine augmentation (0.077–0.150 point improvements) that propagates through the entire workflow. We also saw that memory augmentation closes 33% of the gap between GPT-5-Mini and GPT-5 through improved MITRE mapping, while query construction remains an inherent model capability. Cloud-based multi-step detection remains the most significant challenge (0.282 vs 0.585 for Linux). Future work includes expanding platform coverage and data size, incorporating more tooling systems and exploring different types of agentic harnesses.

### 17 Mar 2026 v2

@@ 4 Experimental Setup
Models Evaluated: We assess 16 model configurations: three Anthropic Claude models (Opus 4.6 at the default high reasoning effort [22], Opus 4.5, Sonnet 4.5), nine OpenAI GPT-5 family variants (GPT-5, 5.1, 5.2 each at high/medium/low reasoning effort), GPT-5-Mini, GPT-4.1, and two reasoning models (O3, O4-Mini). The reasoning effort levels control the computational budget for extended thinking, enabling analysis of how reasoning depth affects detection performance.
@@ 5.1 Overall Model Performance
Claude Opus 4.6 (High) Anthropic 0.6373 0.0374 [0.562, 0.712] 0.266 0.372 0.83 0.50 0.82 0.86 31
Claude Opus 4.5 Anthropic 0.6244 0.0338 [0.556, 0.692] 0.287 0.337 0.91 0.56 0.85 0.92 32
Claude Sonnet 4.5 Anthropic 0.5872 0.0328 [0.521, 0.653] 0.278 0.310 0.85 0.54 0.87 0.88 37
GPT-5 (Med) OpenAI 0.5720 0.0337 [0.504, 0.640] 0.235 0.337 0.72 0.54 0.86 0.38 32
GPT-5.2 (Med) OpenAI 0.5716 0.0325 [0.506, 0.637] 0.226 0.346 0.60 0.37 0.86 0.72 31
GPT-5 (High) OpenAI 0.5633 0.0329 [0.497, 0.629] 0.230 0.333 0.66 0.55 0.86 0.40 32
GPT-5.2 (High) OpenAI 0.5513 0.0367 [0.478, 0.625] 0.232 0.319 0.60 0.40 0.85 0.84 35
GPT-5 (Low) OpenAI 0.5413 0.0315 [0.478, 0.605] 0.204 0.337 0.61 0.45 0.84 0.22 25
GPT-5.1 (Med) OpenAI 0.5133 0.0300 [0.453, 0.574] 0.244 0.269 0.72 0.62 0.87 0.42 26
GPT-5.1 (High) OpenAI 0.4946 0.0373 [0.420, 0.570] 0.238 0.256 0.71 0.55 0.80 0.56 29
GPT-5.2 (Low) OpenAI 0.4898 0.0248 [0.440, 0.540] 0.221 0.269 0.57 0.49 0.88 0.50 27
O3 OpenAI 0.4707 0.0281 [0.414, 0.527] 0.204 0.267 0.51 0.38 0.78 0.68 33
GPT-5-Mini OpenAI 0.4506 0.0292 [0.392, 0.509] 0.188 0.263 0.62 0.21 0.84 0.20 27
GPT-4.1 OpenAI 0.4186 0.0235 [0.371, 0.466] 0.193 0.225 0.68 0.29 0.85 0.02 21
GPT-5.1 (Low) OpenAI 0.3731 0.0211 [0.331, 0.415] 0.194 0.179 0.61 0.37 0.87 0.06 22
O4-Mini OpenAI 0.3602 0.0238 [0.312, 0.408] 0.167 0.193 0.48 0.23 0.80 0.20 26
The results reveal several notable patterns across the 0.637–0.419 performance range. Anthropic Claude models occupy the top three positions, with a clear separation from the best OpenAI result (GPT-5 Med, 0.572). This advantage appears driven by stronger tool-use and agentic capabilities, with particularly large gaps in query execution (C3) and CTI comprehension (C0), as detailed in Section 5.3.
@@ 5.3 Checkpoint Analysis
Figure 4(b) shows normalized mean rewards for each checkpoint. C3 (query execution) is the most discriminating - Claude models score 0.86–0.92 while most OpenAI models fall below 0.50, with GPT-4.1 at just 0.02. C1 (MITRE mapping) varies widely (0.21–0.56), with Claude Opus 4.5 leading at 0.56, suggesting that strong CTI knowledge does not guarantee downstream performance. C0 (CTI report analysis) separates Claude (0.85–0.91) from most OpenAI models (0.49–0.70), reflecting differences in threat intelligence comprehension that propagate through subsequent stages.
@@ 5.6 Ablation: Impact of CTI Tools
Table 3: Ablation study: impact of removing CTI-specific tools. $\Delta$ = Full - Minimal. Per-checkpoint deltas are normalized; C0 excluded (mechanically zeroed without CTI tools).
[…]
Claude Opus 4.6 (High) 0.6130 0.5358 +0.0772 0.30 -0.17 -0.13 -0.12 +0.01
Claude Opus 4.5 0.6124 0.4624 +0.1500 0.59 -0.06 +0.04 +0.03 +0.06
GPT-5 (Med) 0.5454 0.4120 +0.1335 0.64 -0.07 -0.02 +0.16 +0.07
GPT-5.2 (Med) 0.5331 0.4555 +0.0776 0.35 -0.14 -0.03 +0.04 +0.02
GPT-5.1 (Med) 0.4815 0.3641 +0.1174 0.61 -0.05 +0.02 -0.04 +0.04
All five models show performance degradation without CTI tools. Claude Opus 4.5 shows the largest absolute drop ($\Delta=-0.150$), falling from 0.612 to 0.462. GPT-5 (Med) drops by 0.134, and GPT-5.1 (Med) by 0.117. Even the top model, Claude Opus 4.6 (High), drops by 0.077.
Since removing CTI tools mechanically zeroes C0, we exclude it from the per-checkpoint comparison. The right side of Table 3 shows the remaining deltas (positive = full toolkit better). C1 (MITRE mapping) is negative for all models, indicating agents compensate by calling search_mitre_techniques more aggressively without CTI context. C3 (query execution) is positive for three of five models, suggesting agents struggle to construct effective queries without CTI guidance. C4 (detection quality) is non-negative across all models, confirming that the overall reward drop is driven by degraded detection rule quality. CTI context is essential for producing high-quality detections, even if intermediate checkpoints can be reached through alternative paths.
@@ 5.7 Memory Augmentation: Closing the Gap with Larger Models
Memory augmentation closes 33% of the 0.184-point gap to GPT-5 (Med). The improvement is concentrated in knowledge-dependent checkpoints: C1 (MITRE mapping) doubles from 0.22 to 0.44, and the ground truth reward improves by 19%. However, C3 (query execution) remains unchanged at 0.12, suggesting query construction is an inherent model capability rather than a knowledge gap. The remaining gap to GPT-5 lies primarily in C3 (GPT-5 Med achieves 0.40) and detection quality, indicating the larger model’s advantage is in iterative query refinement rather than domain knowledge.
@@
Model Score Checkpoint Reward Ground Truth Reward C0 C1 C2 C3 C4
GPT-5 (Med) 0.5556 0.229 0.326 0.72 0.54 0.80 0.40 0.50
GPT-5 (Low) 0.5478 0.217 0.331 0.64 0.60 0.81 0.20 0.51
GPT-5 (High) 0.5348 0.249 0.286 0.73 0.58 0.83 0.64 0.44
GPT-5-Mini (Memory) 0.4324 0.207 0.225 0.66 0.44 0.86 0.12 0.35
GPT-5-Mini 0.3714 0.182 0.189 0.66 0.22 0.78 0.12 0.29
@@ 8 Discussion and Conclusion
We introduced CTI-REALM, a benchmark for evaluating AI agents on realistic detection engineering workflows. Our evaluation of 16 frontier models reveals that medium reasoning effort consistently outperforms high and low settings. Anthropic Claude models lead overall (Opus 4.6 High: 0.637), and CTI tools provide genuine augmentation (0.019–0.038 point improvements) that propagates through the entire workflow. We also saw that memory augmentation closes 33% of the gap between GPT-5-Mini and GPT-5 through improved MITRE mapping, while query construction remains an inherent model capability. Cloud-based multi-step detection remains the most significant challenge (0.282 vs 0.585 for Linux). Future work includes expanding platform coverage and data size, incorporating more tooling systems and exploring different types of agentic harnesses.